OpenStudy (anonymous):

Will give medels. and become fan!!! Will you exp;ain how to do this type of problem? Solve for x. 6 – 2x – x = 18 A. –24 B. –12 C. –8 D. –4

OpenStudy (twopointinfinity):

subtract 6 from both sides.

OpenStudy (anonymous):

B?

OpenStudy (twopointinfinity):

-2x -3x = x

OpenStudy (twopointinfinity):

divide 12 by -3

OpenStudy (twopointinfinity):

12 / -3 = -4

OpenStudy (anonymous):

-4

OpenStudy (twopointinfinity):

x = -4

OpenStudy (anonymous):

Thank you!!!!! :)

OpenStudy (twopointinfinity):

You're welcome.

OpenStudy (anonymous):

6 – 2x – x = 18 First, 6-(2x+x)=18 Then add 2x and x together, 6-3x=18 Now, -3x=18-6 So, -3x=12 Thus, x=12/-3 x=-4

OpenStudy (twopointinfinity):

Thanks chineseboy. We already answered it, but.. thanks anyway. I'll give you my medal because she gave me hers (:

OpenStudy (anonymous):

Thank you! And your answer is great too!

OpenStudy (anonymous):

Solve. 4x + 2 ≥ –6 A. x ≤ –2 B. x ≤ –1 C. x ≥ –2 D. x ≥ –1

OpenStudy (twopointinfinity):

no, please post it in another question bree (:

OpenStudy (anonymous):

Ok :)